1. A sample of 0.653 g of potassium chloride (KCl) is dissolved in 45.5 g of water. What is the percent by mass of KCl in the solution?


2. Calculate the molality of a sulfuric acid solution containing 34.2 g of sulfuric acid in

173 g of water. The molar mass of sulfuric acid is 98.09 g/mol.


3. How would you prepare a 1.8% NaBr (w:v) aqueous solution?

 


1
Expert's answer
2021-09-16T02:00:37-0400

1)

Percent "=\frac{0.653}{45.5}×100=1.435" %


2)

Moles of sulfuric acid"=\frac{34.2}{98.09}=0.35moles"


Molarity "=\frac{0.35×1000}{173}=2.02M"


3)

To make this solution, you would dissolve 3.3 g

 of sodium bromide in 20.00 mL

 of distilled water, then add enough water to ensure that the final volume of the solution is equal ot 50.00 mL

.